# Optimal Rates for DP-SCO with a Single Epoch and Large Batches

### v1

@@ 4.1 Utility Analysis
Here, we provide the excess population risk guarantee for any choice of the privacy noise random variables $\{\mathbf{b}_{t}\}$’s. We only assume that the variance (over the sampling of the data samples $d\sim\mathcal{D}$) in any individual term in the SRGs (i.e., $\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$) is bounded by $S^{2}$.
Eventually in Theorem 4.6 we will combine 4.1 below with the privacy analysis of Algorithm 1 to arrive at the final DP-SCO guarantee, with the appropriate choice of the batch gradient steps, and the batch size. In particular, there we instantiate the noise random variables $\{\mathbf{b}_{t}\}$’s with the (by now standard) binary tree mechanism [19, 24].
Fix some setting of the $\mathbf{b}_{t}$ such that $\left\|\mathbf{b}_{t}\right\|_{2}\leq b_{max}$ for all $t$. For an appropriate setting of $\eta_{t},\tau_{t}$ and $\beta\geq M$, suppose $\mathbb{E}[\left\|\Delta_{t}-\mathbb{E}[\Delta_{t}]\right\|_{2}^{2}]\leq S^{2}$. Then, over the randomness of the batches $\mathcal{B}$, the following is true for Algorithm 1:
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=$
$O\left(\frac{SL+S^{2}}{\beta\sqrt{T}}+\frac{S\left\|\mathcal{C}\right\|_{2}}{T^{1.5}}+\frac{(S+L)b_{max}}{T}+\frac{\beta b_{max}\left\|\mathcal{C}\right\|_{2}}{T^{2}}+\frac{\beta}{T^{2}}\left\|\mathcal{C}\right\|_{2}^{2}\right).$
We defer the proof to Section 4.4.
Intrpretation of 4.1: To understand the bound in Theorem 4.1, we can view it as three groups of terms. The first group is $\frac{SL+S^{2}}{\beta\sqrt{T}}+\frac{S\left\|\mathcal{C}\right\|_{2}}{T^{1.5}}$, which is the error due to stochasticity of the gradients. The second is $\frac{(S+L)b_{max}}{T}+\frac{\beta b_{max}\left\|\mathcal{C}\right\|_{2}}{T^{2}}$, which is the error due to noise added for privacy. The last is $\frac{\beta}{T^{2}}\left\|\mathcal{C}\right\|_{2}^{2}$, which is the optimization error, i.e. the error of the algorithm without stochasticity or DP noise. Notice that the optimization error matches that of the non-stochastic part of non-private Nesterov’s acceleration [41]. Without privacy, the stochastic part of Nesterov’s acceleration depends on $\frac{V}{\beta\sqrt{T}}$ [41], where $V$ bounds stochasticity in evaluating the population gradient: $\mathbb{E}_{\mathcal{B}_{t}}\left[\left\|\frac{1}{|\mathcal{B}_{t}|}\sum\limits_{d\in\mathcal{B}_{t}}\nabla f(\mathbf{x};d)-\nabla F(\mathbf{x};d)\right\|_{2}^{2}\right]\leq V^{2}$. Morally, the parameter $S$ in 4.1 can be compared to $V$. However, $S$ encodes the additional variance due to stochastic recursive gradients. Modulo, the difference between $S$ and $V$4.1 cleanly decomposes the non-private stochastic Nesterov’s utility and the additional error due to the explicit noise introduced due to DP.
Proof overview: The proof is deferred to Section 4.4. We give a high-level overview here. We define the following potential. Let $\eta_{0:t}=\sum\limits_{i=0}^{t}\eta_{i}$, then
Independent batch gradients: In Theorem 4.1 we analyzed Algorithm 1 where the $\nabla_{t}$ are computed using SRG. The analysis can easily be extended to the standard setting where each $\nabla_{t}$ is an independent batch gradient. We give this analysis in Section 4.4.3. To the best of our knowledge, a direct potential-based analysis of stochastic accelerated gradient descent has not appeared in the literature before besides [51]. Our analysis is simpler than that of [51], so it may be of independent interest.
4.2 Privacy analysis
Given Theorem 4.1, our goal now is to bound $S$, which is equivalent to bounding the sensitivity of the $\Delta_{t}$, and then to show that there is a noise distribution for which $b_{max}$ is sufficiently small (with high probability) and we achieve $(\varepsilon,\delta)$-DP. We can then set $\beta$ to minimize the error bound, and determine the largest batch size that doesn’t cause the error bound to increase.
We first prove the sensitivity bound on each $\Delta_{t}$. The sensitivity bound requires a bound on the max noise we add in any iteration. The proof of sensitivity follows a similar structure to that of [55, Lemma 15]. However, because the gradients $\nabla f_{t}(\mathbf{x}_{t})$ are evaluated at a different point than that in [55], the exact argument is different.
Claim 4.2 (Bounding $\ell_{2}$-sensitivity).
$\left\|\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)\right\|_{2}\leq 2(Mb_{\max}+2M\left\|\mathcal{C}\right\|_{2}+L).$
Proof.
$\leq\eta_{0:t-1}\left\|\frac{\nabla_{t-1}}{\beta}+\frac{\mathbf{b}_{t-1}}{\eta_{t}}\right\|_{2}+\eta_{t}\left\|\mathcal{C}\right\|_{2}$
$\Rightarrow\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$ $\leq\frac{\eta_{0:t-1}}{\eta_{0:t}}\cdot\frac{\left\|\nabla_{t-1}\right\|_{2}}{\beta}+\frac{\eta_{0:t-1}\left\|\mathbf{b}_{t-1}\right\|_{2}}{\eta_{t}\cdot\eta_{0:t}}+\frac{\eta_{t}}{\eta_{0:t}}\left\|\mathcal{C}\right\|_{2}$
$\leq\frac{\sum_{i\leq t-1}\sum_{d\in\mathcal{B}_{i}}\left\|\Delta_{i}(d)\right\|_{2}}{\beta(t+1)B}+\frac{1}{t+1}b_{\max}+\frac{2}{t+2}\left\|\mathcal{C}\right\|_{2}$
$\leq\frac{\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{\beta}+\frac{1}{t+1}b_{\max}+\frac{2}{t+2}\left\|\mathcal{C}\right\|_{2}$
$\Rightarrow\eta_{t}\cdot M\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$ $\leq\frac{\eta_{t}M\cdot\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{\beta}+Mb_{\max}+2M\left\|\mathcal{C}\right\|_{2}$
$\leq\frac{\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{2}+Mb_{\max}+2M\left\|\mathcal{C}\right\|_{2}.$ (2)
$\left\|\Delta_{t}(d)\right\|_{2}\leq\frac{\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{2}+Mb_{max}+2M\left\|\mathcal{C}\right\|_{2}+L.$ (3)
By induction, this completes the proof of 4.2. ∎
A crucial aspect of 4.2 is that conditioned on $b_{max}$, the sensitivity is independent of $t$, even if the norm of the quantity that is being approximated (i.e., $\eta_{t}\nabla f(\mathbf{x}_{t};d)$) scales with $t$.
Lemma 4.3.
Let $\mathbf{g}_{i}=\sum_{j\leq i}\Delta_{i}$ for $i\in[T]$ and suppose the $\Delta_{i}$ are computed such that between any two adjacent datasets, only one $\Delta_{i}$ changes, and it changes by $\ell_{2}$-norm at most $C$. Then for any $T\geq 2$ there is a $\mu$-GDP algorithm (i.e., an algorithm which satisfies any DP guarantee satisfied by the Gaussian mechanism with sensitivity $\mu$ and variance 1) that gives an unbiased estimate of each $\mathbf{g}_{i}$ such that with probability $1-\delta$ all $\mathbf{g}_{i}$’s estimates have $\ell_{2}$-error at most $\frac{4C\log^{1.5}T\sqrt{d\ln(2T/\delta)}}{\mu}$.
Proof.
Each $\Delta_{i}$ contributes to at most one $s_{j,k}$ for each value of $k$, i.e. at most a total of $1+\lceil\log T\rceil$ values of $s_{j,k}$. So, if $\left\|\Delta_{i}\right\|_{2}\leq C$ for all $i$, then with $\sigma=C\sqrt{1+\lceil\log T\rceil}/\mu$ we can release all $s_{j,k}$ with $\mu$-GDP. Furthermore, by a standard Gaussian tail bound with probability $21-\delta/$ 2 we have that $\left\|\xi_{i}\right\|_{2}\leq\sigma\sqrt{2d\ln(2T/\delta)}$ for all $i$. Since each $\mathbf{g}_{i}$ has an unbiased estimator formed by adding at most $\lceil\log T\rceil$ of the $s_{j,k}$, the maximum additive error is then at most
$\lceil\log T\rceil\max_{i}\left\|\xi_{i}\right\|_{2}\leq 2C\log T\sqrt{2d(1+\lceil\log T\rceil)\ln(2T/\delta)}/\mu\leq 4C\log^{1.5}T\sqrt{d\ln(2T/\delta)}/\mu.$
4.3 gives us a high probability bound on the max noise, which we can then plug into 4.2 to show that if the dimension (and thus the norm of the noise) is not too large, the noise term in 4.2 is not dominating:
Corollary 4.4.
Suppose we compute $\mathbf{b}_{i}$ in Algorithm 1 using the $\mu$-GDP binary tree mechanism (Lemma 4.3) with $C=\frac{4L+8M\left\|\mathcal{C}\right\|_{2}}{B\beta}$. Then for $d\leq\frac{B^{2}\beta^{2}\mu^{2}}{64M^{2}\log^{3}(T)\ln(4T/\delta)}$ and the parameter choices in Claim 4.2, with probability $21-\delta/$ 2, each $\Delta_{t}$ has $\ell_{2}$-sensitivity at most $\leq 4L+8M\left\|\mathcal{C}\right\|_{2}$ for all $i$.
Our eventual choice of $\beta$ will be proportional to $n^{3/2}/B^{2}$, i.e. the upper bound on $d$ is proportional to $n^{3}\varepsilon^{2}/B^{2}$. We can rewrite this as an upper bound on the batch size: $B\leq n^{1.5}\varepsilon/\sqrt{d}$. Ignoring dependence on logarithmic terms and $\delta$, when $d\geq n^{2}\varepsilon^{2}$, no $(\varepsilon,\delta)$-DP algorithm can achieve expected excess loss, $o(L\left\|\mathcal{C}\right\|_{2})$, which is achieved by any point in $\mathcal{C}$. So this upper bound is vacuous in the setting where $d\leq n$ and $\varepsilon=\Omega(1)$, and never worse than $\sqrt{n}$ for interesting settings of $\delta$.
We can now convert the sensitivity bound and the privacy guarantees of the binary tree mechanism into a privacy guarantee for Algorithm 1:
Corollary 4.5.
Suppose we compute $\mathbf{b}_{i}$ in Algorithm 1 using the binary tree mechanism (Lemma 4.3) with $\sigma=\frac{(8\sqrt{2}L+16\sqrt{2}M\left\|\mathcal{C}\right\|_{2})\sqrt{\log(T)\ln(2.5/\delta)}}{\varepsilon B\beta}$ on $\Delta_{i}$. Then for $d\leq\frac{B^{2}\beta^{2}\varepsilon^{2}}{128M^{2}\log^{3}(T)\ln(4T/\delta)\ln(2.5/\delta)}$ and the parameter choices in Claim 4.2, Algorithm 1 satisfies $(\varepsilon,\delta)$-DP.
Proof.
If we run Algorithm 1 but instead clip $\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$ for each $t$ and $d$ in $\mathcal{B}_{t}$ to norm at most $4L+8M\left\|\mathcal{C}\right\|_{2}$ if its norm exceeds $4L+8M\left\|\mathcal{C}\right\|_{2}$, still using the binary tree mechanism to generate the $\mathbf{b}_{i}$, then Algorithm 1 satisfies $\varepsilon/\sqrt{2\ln(2.5/\delta)}$-GDP which implies it satisfies $(\varepsilon,\delta/2)$-DP. Algorithm 1 without clipping is within total variation distance $\delta/2$ of the version with clipping by Corollary 4.4, i.e. Algorithm 1 satisfies $(\varepsilon,\delta$)-DP. ∎
Theorem 4.6.
[Simplification of Theorem 4.10 in Section 4.4] Algorithm 1 with batch size $B=\sqrt{n}\cdot\min\{1,M\}$ and $T=n/B$ (i.e., a single epoch over a dataset of size $n$), and an appropriate choice of $\eta_{t},\beta,\sigma$ is $(\varepsilon,\delta)$-DP and has the following excess population loss guarantee:
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=O\left((L\left\|\mathcal{C}\right\|_{2}+M\left\|\mathcal{C}\right\|_{2}^{2})\cdot\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{d}}{\varepsilon n}\cdot\text{polylog}(n/B,1/\delta)\right)\right).$
4.3 Unaccelerated Variant
A natural question is whether an unaccelerated variant of Accelerated-DP-SRGD can also achieve optimal DP-SCO rates with batch size $\omega(1)$. We give a sketch that the unaccelerated version requires batch size $O(1)$. Suppose we use $\Delta_{t}=c_{t}\nabla f(\mathbf{x}_{t},\mathcal{B}_{t})-c_{t-1}\nabla f(\mathbf{x}_{t-1},\mathcal{B}_{t})$ as our gradient differences and $\nabla_{t}=\frac{\sum_{i\leq t}\Delta_{i}}{c_{t}}$ as our stochastic recursive gradients. For simplicity we will consider using SGD with a fixed learning rate $\eta$ as the underlying optimization algorithm, assume $L=\left\|\mathcal{C}\right\|_{2}=M=1$ and assume $c_{t}$ grows at most polynomially, which implies $\sum_{i\leq t}c_{i}^{2}\approx tc_{t}^{2}$. Similarly to Claim 4.2, we can show the sensitivity of each $\Delta_{t}$ is roughly $c_{t}\eta$. Suppose we use $\eta\approx\sqrt{B/T}=B/\sqrt{n}$, which is standard for interpolation between full-batch GD and SGD with $B=1$ (see e.g. Section 6 of [12]). We compute the vector variance $\mathbf{Var}\left(\mathbf{v}\right):=\mathbb{E}[\left\|\mathbf{v}-\mathbb{E}[\mathbf{v}]\right\|_{2}^{2}]$ of $\nabla_{t}$:
$\mathbf{Var}\left(\nabla_{t}\right)=\mathbf{Var}\left(\frac{\sum_{i\leq t}\Delta_{t}}{c_{t}}\right)=\frac{\sum_{i\leq t}\mathbf{Var}\left(\Delta_{t}\right)}{c_{t}^{2}}=\frac{\sum_{i\leq t}c_{i}^{2}\eta^{2}}{c_{t}^{2}}=\frac{t\eta^{2}}{B}=\frac{tB}{n}.$
4.4 Deferred Proofs
4.4.1 Proof of Theorem 4.1
Theorem 4.1 is given by plugging $\eta_{t}={t+1}$ into the following more general theorem, dropping lower-order terms, setting $S=Q/\sqrt{B}$, and taking the asymptotic bound.
@@ 4.4 Deferred Proofs
Then, over the randomness of the i.i.d. sampling of the loss functions $\{f(\cdot;d_{t})\}$ from $\mathcal{D}$, the following is true for Algorithm 1:
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})\leq$
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}+\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+\frac{Q\left\|\mathcal{C}\right\|_{2}Q(t+1)}{\sqrt{B}\eta_{0:t}}+\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)$
$+\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}.$
@@
Let $\mathbf{q}_{t}=\nabla_{t}-\nabla F(\mathbf{x}_{t})$. Assuming $\eta_{t}^{2}\leq 4\eta_{0:t}$, the potential drop is bounded by the following (where $\Delta\Phi_{t}=\Phi_{t+1}-\Phi_{t}$):
$\mathbb{E}\left[\Delta\Phi_{t}\right]\leq\eta_{0:t}\left(\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}+\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{t+1}}{\sqrt{B}\eta_{0:t}}\right.$
Telescopic summation and averaging: By telescoping the potential differences $\Delta\Phi_{t}$’s and linearity of expectation we have the following over the sampling of the functions $\{f_{t}\}$’s:
$\mathbb{E}[\Phi_{T}-\Phi_{0}]\leq$ $\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}+\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{t+1}}{\sqrt{B}\eta_{0:t}}\right.$
@@
Therefore, from Equation 5 and definition of $\Phi_{t}$, as desired we have:
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})\leq$
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}+\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{t+1}}{\sqrt{B}\eta_{0:t}}\right.$
$\left.+\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)+\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}.$
@@
Using 4.9 below, and the fact that $\tau_{t}\leq 1$, we have the potential drop in 4.8 bound by the following.
@@
Recall, by definition $\mathbf{y}_{t+1}=\operatorname*{arg\,min}\limits_{\mathbf{y}\in\mathcal{C}}\left\langle\mathbf{q}_{t}+\nabla F(\mathbf{x}_{t})-\frac{\beta}{\eta_{t}}\cdot\mathbf{b}_{t},\mathbf{y}-\mathbf{x}_{t}\right\rangle+\frac{\beta}{2}\left\|\mathbf{y}-\mathbf{x}_{t}\right\|_{2}^{2}$. Let $\mathbf{y}_{t+1}^{\prime}=\mathbf{x}_{t}-\frac{1}{\beta}\nabla_{t}+\frac{1}{\eta_{t}}\cdot\mathbf{b}_{t}$ be the unconstrained minimizer, and let $J(\mathbf{y})=\left\langle\mathbf{q}_{t}+\nabla F(\mathbf{x}_{t})-\frac{\beta}{\eta_{t}}\cdot\mathbf{b}_{t},\mathbf{y}\right\rangle+\frac{\beta}{2}\left\|\mathbf{y}\right\|_{2}^{2}$. By definition, $J(\mathbf{y}_{t+1})\geq J(\mathbf{y}^{\prime}_{t+1})+\frac{\beta}{2}\left\|\mathbf{y}_{t+1}-\mathbf{y}^{\prime}_{t+1}\right\|_{2}^{2}$. Therefore,

### v2

@@ 5.1 Utility Analysis
First, we provide the excess population risk guarantee for any choice of the privacy noise random variables $\{\mathbf{b}_{t}\}$’s. We only assume that the variance (over the sampling of the data samples $d\sim\mathcal{D}$) in any individual term in the SRGs (i.e., $\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$) is bounded.
@@
Then, w.p. $1-\gamma$ over the randomness of the i.i.d. sampling of the batches $\{\mathcal{B}_{t}\}$, the following is true for Algorithm 1:
$F(\mathbf{y}_{T})-F(\mathbf{x}^{*})\leq$
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{3(t+1)\log(6T/\gamma)}(\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}+\frac{Q\sqrt{3(t+1)\log(6T/\gamma)}}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}\right.$
$\left.+\frac{3Q\sqrt{3(t+1)\log(6T/\gamma)}\left\|\mathbf{b}_{t}\right\|_{2}}{\sqrt{B}\eta_{t}^{2}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{3(t+1)\log(6T/\gamma)}}{\sqrt{B}\eta_{0:t}}+\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}\right.$
$\left.+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)+\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}.$
@@
Let $\mathbf{q}_{t}=\nabla_{t}-\nabla F(\mathbf{x}_{t})$. Assuming $\eta_{t}^{2}\leq 4\eta_{0:t}$, w.p. $1-\gamma$ over $\{\mathcal{B}_{t}\}$ the potential drop is bounded by the following (where $\Delta\Phi_{t}=\Phi_{t+1}-\Phi_{t}$):
$\Delta\Phi_{t}$ $\leq\eta_{0:t}\left(\frac{2Q\sqrt{3(t+1)\log(6/\gamma)}(\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}+\frac{Q\sqrt{3(t+1)\log(6/\gamma)}}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}\right.$
$+\frac{3Q\sqrt{3(t+1)\log(6/\gamma)}\left\|\mathbf{b}_{t}\right\|_{2}}{\sqrt{B}\eta_{t}^{2}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{3(t+1)\log(6/\gamma)}}{\sqrt{B}\eta_{0:t}}$
$+\left.\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right).$
Telescopic summation and averaging: By telescoping the potential differences $\Delta\Phi_{t}$’s and a union bound over failure probabilities we have the following w.p. $1-\gamma$ over the sampling of the functions $\{f_{t}\}$’s:
$\Phi_{T}-\Phi_{0}\leq$ $\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{3(t+1)\log(6T/\gamma)}(\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}+\frac{Q\sqrt{3(t+1)\log(6T/\gamma)}}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}\right.$
$+\frac{3Q\sqrt{3(t+1)\log(6T/\gamma)}\left\|\mathbf{b}_{t}\right\|_{2}}{\sqrt{B}\eta_{t}^{2}}\left.+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{3(t+1)\log(6T/\gamma)}}{\sqrt{B}\eta_{0:t}}\right.$ (2)
@@
Therefore, from Equation 3 and definition of $\Phi_{t}$, as desired we have w.p. $1-\gamma$:
$F(\mathbf{y}_{T})-F(\mathbf{x}^{*})\leq$
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{3(t+1)\log(6T/\gamma)}(\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}+\frac{Q\sqrt{3(t+1)\log(6T/\gamma)}}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}\right.$
$+\frac{3Q\sqrt{3(t+1)\log(6T/\gamma)}\left\|\mathbf{b}_{t}\right\|_{2}}{\sqrt{B}\eta_{t}^{2}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{3(t+1)\log(6T/\gamma)}}{\sqrt{B}\eta_{0:t}}+\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}$
$\left.+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)+\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}.$
@@
Using Lemma 5.3 below, and the fact that $\tau_{t}\leq 1$, we have the potential drop in Lemma 5.2 bound by the following.
@@
Recall, by definition $\mathbf{y}_{t+1}=\operatorname*{arg\,min}\limits_{\mathbf{y}\in\mathcal{C}}\left\langle\mathbf{q}_{t}+\nabla F(\mathbf{x}_{t})-\frac{\beta}{\eta_{t}}\cdot\mathbf{b}_{t},\mathbf{y}-\mathbf{x}_{t}\right\rangle+\frac{\beta}{2}\left\|\mathbf{y}-\mathbf{x}_{t}\right\|_{2}^{2}$. Let $\mathbf{y}_{t+1}^{\prime}=\mathbf{x}_{t}-\frac{1}{\beta}\nabla_{t}+\frac{1}{\eta_{t}}\cdot\mathbf{b}_{t}$ be the unconstrained minimizer, and let
$J(\mathbf{y})=\left\langle\mathbf{q}_{t}+\nabla F(\mathbf{x}_{t})-\frac{\beta}{\eta_{t}}\cdot\mathbf{b}_{t},\mathbf{y}\right\rangle+\frac{\beta}{2}\left\|\mathbf{y}\right\|_{2}^{2}.$
By definition, $J(\mathbf{y}_{t+1})\geq J(\mathbf{y}^{\prime}_{t+1})+\frac{\beta}{2}\left\|\mathbf{y}_{t+1}-\mathbf{y}^{\prime}_{t+1}\right\|_{2}^{2}$. Therefore,
@@
$+3\beta\left\|\mathbf{b}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}+\eta_{t}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}$
$\leq$ $\frac{3\eta_{0:t}}{\beta}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\nabla_{t}\right\|_{2}+\frac{3\eta_{0:t}}{\eta_{t}}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\mathbf{b}_{t}\right\|_{2}+\frac{L\left\|\mathbf{b}_{t}\right\|_{2}}{\tau_{t}}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}\tau_{t}}$